# Electromagnetic Follow-up to Gravitational Wave Events with the UltraViolet EXplorer (UVEX)

### v2

@@ 1.1 UVEX: The UltraViolet EXplorer
As of February 13th, 2024, the UltraViolet EXplorer (UVEX; PI Fiona Harrison) has been selected by NASA as its next Astrophysics Medium-Class Explorer mission with a launch date of 2030 (Fisher, 2024). UVEX will be a space telescope with wide-field, co-aligned imaging in the near- and far-ultraviolet (NUV/FUV) band with a 12.25 deg2 field of view. It will additionally possess moderate-resolution spectroscopic capabilities across a broad bandpass covering much of the FUV and NUV (Kulkarni et al., 2023). In addition to the rapid target-of-opportunity (ToO) capabilities considered in this work, UVEX will serve as a joint FUV/NUV deep survey instrument, providing a crucial multi-band UV counterpart to the optical and infrared surveys of the coming era (Rubin, Euclid, and Roman; Ivezić et al., 2019; Laureijs et al., 2011; Akeson et al., 2019). These capabilities allow UVEX to cover a missing sector of the observational landscape of the 2030s: a deep, multi-band UV mission with ample sky coverage and rapid ToO response times. The Hubble Space Telescope, while of course highly sensitive, possesses a narrow field of view and slow ($>2$ weeks) ToO response times; the Neil Gehrels Swift Observatory (Swift; Gehrels et al., 2004) only covers the NUV; and the Ultraviolet Transient Astronomical Satellite (ULTRASAT; Sagiv et al., 2014; Shvartzvald et al., 2024), while effective for rapid ToOs and imaging in the NUV, does not provide the depth, spectral capabilities, or FUV coverage of UVEX. For further details as to the planned capabilities and broad science portfolio of UVEX, including its planned surveys and time-domain studies beyond those considered here, refer to Kulkarni et al. (2023).
@@ 1.3 Challenges for Electromagnetic Follow-up
Numerous thorough searches were performed for the BNS merger GW190425 (Abbott et al., 2020a) (e.g., Antier et al., 2020a; Coughlin et al., 2019b; Gompertz et al., 2020; Hosseinzadeh et al., 2019; Saleem et al., 2020; Song et al., 2019; Paek et al., 2024), as well as the NSBH mergers GW200105 and GW200115 (Abbott et al., 2021a) (e.g., Page et al., 2020; Antier et al., 2020b; Anand et al., 2021; Dichiara et al., 2021) and GW events in O4 (Ahumada et al., 2024). Part of the challenge lies in the broad sky area that any such follow-up endeavor must search. GW detectors do not have high angular resolution to individual sources, and while a network of 3+ detectors allows for significant improvements in this respect, GW localization regions — that is, the area of the sky covered by, e.g., the 90% credible level (C.L.) error region for the GW-inferred localization — are expected to remain large (many 10s or 100s of square degrees) for the immediate future (Petrov et al., 2022; Kiendrebeogo et al., 2023). Tiling a GW localization region spanning such a broad area of the sky is a difficult task for most modern observatories. Adding this difficulty, light curve predictions from KN models for these events are still quite uncertain. This is largely due to the fact that we do not yet understand the dominant mechanism powering the KN at early times. Possibilities include a predominantly nucleosynthesis-powered KN (Li & Paczyński, 1998; Metzger et al., 2010; Hotokezaka & Nakar, 2020; Banerjee et al., 2020; Klion et al., 2021; Banerjee et al., 2024), a nucleosynthesis-powered KN with additional contributions from $\beta$-decay of free neutrons (Metzger et al., 2015; Gottlieb & Loeb, 2020), and a shock-powered KN (Piro & Kollmeier, 2018). Depending on which of these prescriptions is used, the corresponding light-curve predictions can vary by more than 4 mag at early times (see, e.g., Fig. 43 of Kulkarni et al., 2023). When this uncertainty is joined by the intrinsic GW uncertainties on the inclination $\iota$ and luminosity distance $D_{L}$ of the BNS/NSBH merger in question, the depth at which a given telescope needs to observe each field to be able to actually detect a potential KN becomes difficult to estimate. One could, of course, take a highly conservative approach and observe each field at extreme depth, but doing so requires large per-field exposure times.
@@
Time is, however, of the essence. Not only are KNe fast-evolving and fade quickly (as seen for AT 2017gfo in, e.g., Cowperthwaite et al., 2017), but many open questions as to the physical nature of KNe and the origin of heavy elements can only be answered by observing the KN during the rapid evolution of its light curve at early times ($\lesssim 6$ hrs in the UV; see discussion in Appendix E of Kulkarni et al., 2023). Successful EM-GW follow-up searches must therefore be deep, broad, and rapid. This quandary is the driving motivation for employing wide-field survey telescopes in the pursuit of EM follow-up to GW events. These observatories are designed with wide fields of view and the ability to quickly move from pointing to pointing for efficient survey operations. The primary challenge then becomes how one can achieve the necessary depth for successful followup of a given KN counterpart in the short time before it fades without sacrificing coverage of the GW localization region.
@@ 4 Estimated Success Rates
Additional complicating factors (extinction, UV background, etc.) are accounted for within the UVEX ETC. The true apparent magnitude of the counterpart, $m_{\mathrm{KN}}$, is given in terms of the true intrinsic KN absolute magnitude and the actual source distance:
@@
The prior bounds are chosen so as to encapsulate the full extent of KN FUV emission models considered in Kulkarni et al. (2023); see Fig. 43 and surrounding discussion therein. We can now write an expression for the probability of success in terms of depth, $p_{\mathrm{depth}}(\mathcal{S})$:
@@ 5 Updated Prospects for UVEX EM-GW Follow-up
The prospects presented here for UVEX EM-GW follow-up build on the procedure developed for the UVEX proposal and selection process, bringing it in line with the current post-selection UVEX timeline and mission details — although these remain subject to change — as well as our improved understanding of the BNS rate and mass distribution. As the precise details of the LVK observing run timeline and detector sensitivity in the coming years remains uncertain, we consider two fiducial observing scenarios. The “O5” scenario considers 1.5 years of overlap with a LVK detector network consisting of the LIGO Hanford, LIGO Livingston, Virgo, and KAGRA at their respective projected O5 sensitivities. The “O6” scenario considers the same configuration at the projected sensitivities for O6 and a fiducial 1.5 year overlap. These prospects are estimated by applying our UVEX EM-GW ToO selection criteria as developed for the UVEX proposal to the updated scenarios of Kiendrebeogo et al. (2023) with the GWTC-3 astrophysical BNS rate estimate of $210^{+240}_{-120}\ \mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$ (Abbott et al., 2023). The uvex-followup configuration used for both the O5 and O6 simulations is given in Table 2. These updated estimates retain the UVEX ToO selection strategy developed for the UVEX proposal, and can therefore be taken as a modernization of the quotes presented in Kulkarni et al. (2023).888i.e., 20 ToO triggers for O5 and 35 for O6; the (Kulkarni et al., 2023) quotes did not include the detailed treatment of uncertainty now present within uvex-followup (as described in §3.3). We consider an expanded set of ToO selection strategies in §6.
For the O5 (O6) simulation, with the fiducial observing strategy detailed in Table 2 and an overlap of 1.5 years of concurrent observations, UVEX is estimated to trigger EM-GW ToO observations for $8.9^{+11.1}_{-5.0}$ ($11.3^{+14.0}_{-6.3}$) BNS mergers. The success rate of the fiducial observing strategy as applied to the O5 (O6) scenarios is 76.3% (77.5%), corresponding to an estimated $6.8^{+8.4}_{-3.8}$ ($8.8^{+10.9}_{-4.9}$) KN counterpart detections. The overall distribution of these ToO triggers with respect to their GW distance estimates and sky localizations for O5 and O6 can be found in Figs. 1 and 2, respectively. It is expected that the LVK BNS rate measurement will change following O4; to allow the estimates presented in this work to be robust to rate updates in the coming years, we provide a simple tool to convert these quotes to the corresponding quote with a BNS rate of the user’s choice. This tool can be found at https://github.com/criswellalexander/uvex-rate-updater.
@@
Figure 1: Fiducial UVEX ToO selection for the O5 simulation. Points correspond to individual simulated events from the Kiendrebeogo et al. (2023) observing scenario, so the total number of points is not one-to-one with the expected number of observed BNS mergers (although it is proportional); scaling from the simulated event rate to an astrophysical BNS rate of $210^{+240}_{-120}\ \mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$ (Abbott et al., 2023) yields $8.9^{+11.1}_{-5.0}$ UVEX ToO triggers over 1.5 years of O5. Events are plotted by their mean GW distance estimate and 90% C.L. GW localization areas. Points in grey are not selected due to the overall threshold of $\leq 100$ deg2 GW 90% C.L. localization area. Events with $\leq 100$ deg2 localizations that are not selected due falling outside of the UVEX FoR or exceeding the maximum epoch time constraint of 3 hrs are marked as black and orange $\times$×s, respectively. Selected events are color-coded by per-tile exposure time.
Figure 2: UVEX ToO selection for the O6 simulation. Points correspond to individual simulated events from the O6 bns_broad observing scenario, so the total number of points is not one-to-one with the expected number of observed BNS mergers (although it is proportional); scaling from the simulated event rate to an astrophysical BNS rate of $210^{+240}_{-120}\ \mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$ (Abbott et al., 2023) yields $11.3^{+14.0}_{-6.3}$ UVEX ToO triggers over 1.5 years of O6. Events are plotted by their mean GW distance estimate and 90% C.L. GW localization areas. Points in grey are not selected due to the overall threshold of $\leq 100$ deg2 GW 90% C.L. localization area. Events with $\leq 100$ deg2 localizations that are not selected due falling outside of the UVEX FoR or exceeding the maximum epoch time constraint of 3 hrs are marked as black and orange $\times$×s, respectively. Selected events are color-coded by per-tile exposure time.
@@ 6.2 Coverage Threshold
The percent coverage threshold defines the minimum allowed coverage by UVEX of the GW localization region. It is given as the total integrated probability tiled by the observing plan; e.g., 99% coverage equates to a 99% chance that the counterpart is present within the observed fields. Higher coverage thresholds yield a higher confidence that the UVEX ToO observing plan will include the KN counterpart; however, a stringent threshold may reject events with a high probability of success, but which are missing some small area of the localization region due to sun exclusion, time constraints, etc.. It is worth noting that the threshold defines a minimum level of coverage, and therefore most selected events will in fact encapsulate a greater total probability than given by the threshold. The (quite conservative) fiducial coverage threshold employed for the UVEX proposal process was 99%. We consider two variant strategies — dubbed CT95 and CT90 — in which this threshold is relaxed to 95% and 90%, respectively, and consider the impact on the overall ToO count and success rate. Doing so yields a small increase in ToO triggers with a negligible change in success rate: adopting the CT90 strategy leads to $3.2^{+3.9}_{-1.8}$ and $3.4^{+4.3}_{-2.0}$ additional triggers for the O5 and O6 scenarios, for a total of $12.1^{+15.0}_{-6.8}$ and $14.7^{+18.3}_{-8.3}$ triggers, respectively. The success rate of the CT90 strategy as applied to the O5 scenario is 0.3% higher than that of the fiducial case and is only reduced by 0.2% for the O6 scenario, yielding a total of $9.3^{+11.5}_{-5.2}$ ($11.4^{+14.1}_{-6.4}$) detected counterparts for O5 (O6). The relative ToO trigger counts and success rates for coverage thresholds of 99%, 95%, and 90% can be found in Table 3. An event selection plot for the CT90 strategy can be found in Fig. 3. In general, relaxing this requirement results better retention of nearby/well-localized events that fall marginally outside of the UVEX FoR; compare, e.g., the proportion of events lost to FoR exclusion in Fig. 2 to that of Fig. 3.
Figure 3: UVEX ToO selection for the O6 simulation with a 90% threshold for minimum coverage by UVEX of the GW localization. Points correspond to individual simulated events from the Kiendrebeogo et al. (2023) O6 observing scenario, so the total number of points is not one-to-one with the expected number of observed BNS mergers (although it is proportional); scaling from the simulated event rate to an astrophysical BNS rate of $210^{+240}_{-120}\ \mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$ (Abbott et al., 2023) yields $14.7^{+18.3}_{-8.3}$ UVEX ToO triggers over 1.5 years of O6. Events are plotted by their mean GW distance estimate and 90% C.L. GW localization areas. Points in grey are not selected due to the overall threshold of $\leq 100$ deg2 GW 90% C.L. localization area. Events with $\leq 100$ deg2 localizations that are not selected due falling outside of the UVEX FoR or exceeding the maximum epoch time constraint of 3 hrs are marked as black and orange $\times$×s, respectively. Selected events are color-coded by per-tile exposure time.

### v4

@@ 1.1 UVEX: The UltraViolet EXplorer
As of February 13th, 2024, the UltraViolet EXplorer (UVEX; PI Fiona Harrison) has been selected by NASA as its next Astrophysics Medium-Class Explorer mission with a launch date of 2030 (Fisher, 2024). UVEX will be a space telescope with wide-field, co-aligned imaging in the near- and far-ultraviolet (NUV/FUV) band with a 12.25 deg2 field of view. It will additionally possess moderate-resolution ($R\geq 1000$) spectroscopic capabilities across a broad bandpass covering much of the FUV and NUV (Kulkarni et al., 2023). In addition to the rapid target-of-opportunity (ToO) capabilities considered in this work, UVEX will serve as a joint FUV/NUV deep survey instrument, providing a crucial multi-band UV counterpart to the optical and infrared surveys of the coming era (Rubin, Euclid, and Roman; Ivezić et al., 2019; Laureijs et al., 2011; Akeson et al., 2019). These capabilities allow UVEX to cover a missing sector of the observational landscape of the 2030s: a deep, multi-band UV mission with ample sky coverage and rapid ToO response times. The Hubble Space Telescope, while of course highly sensitive, possesses a narrow field of view (0.002 deg2) and slow ($>2$ weeks) ToO response times; the Neil Gehrels Swift Observatory (Swift; Gehrels et al., 2004) only covers the NUV; and the Ultraviolet Transient Astronomical Satellite (ULTRASAT; Sagiv et al., 2014; Shvartzvald et al., 2024), while effective for rapid ToOs and imaging in the NUV with a planned ToO response time of $<15$ min from ToO trigger and a FOV of 204 deg2, does not provide the depth, spectral capabilities, or FUV coverage of UVEX. For detailed comparison of the relevant capabilities of current and planned UV observatories to UVEX and typical requirements for follow-up to gravitational wave (GW) events, see Table 1. For further details as to the planned capabilities and broad science portfolio of UVEX, including its planned surveys and time-domain studies beyond those considered here, refer to Kulkarni et al. (2023).
Swift Hubble ULTRASAT UVEX O6 BNS/KNe
UV Band(s) NUV NUV NUV NUV+FUV NUV+FUV
FOV 0.08 deg2 0.002 deg2 204 deg2 12.25 deg2 10s – 100s of deg2
Depth (AB mag) 24 28.75 22.5 24.7 (NUV)/24.6 (FUV) 19.6–25.0 (NUV at 300 Mpc)
ToO Response Time $<150$ s $>2$ weeks $<15$ min $3$ hrs (average) $\sim$6 hrs
Table 1: UVEX capabilities compared to other current and upcoming UV observatories and typical expected values for LVK-detected BNS mergers and their KN counterparts in O6. The “FOV” values given for O6 BNS/KNe are the expected sky localizations for well-localized BNS in O6 (Petrov et al., 2022). The range of KN apparent magnitudes corresponds to the minimum/maximum peak NUV luminosity for KN emission models considered in Kulkarni et al. (2023), observed at a distance of 300 Mpc (the equivalent quote for the fainter FUV emission is 19.5–27.2 AB mags). The ToO response time listed for BNS mergers is the expected timescale after which it will likely no longer be possible distinguish between emission models for the KN UV emission; see e.g. Figure 14 of Kulkarni et al. (2023). The depths quoted for ULTRASAT and UVEX are limiting magnitudes for a $5\sigma$ detection in a standard 900 s dwell. The depth quoted for Hubble is the current limiting magnitude of WFC3/UVIS at 2500Å as given by Figure 3 of the Hubble Space Telescope Primer for Cycle 33 (Peeples & The STScI Science Mission Office, 2025). The depth quoted for Swift is the $5\sigma$ limiting magnitude of the UVOT instrument for a 1000 s exposure with the white light filter (Gehrels et al., 2004). The ToO response time for Swift is based on updated capabilities presented in (Tohuvavohu et al., 2024). Remaining quotes drawn from Gehrels et al. (2004), Peeples & The STScI Science Mission Office (2025), Shvartzvald et al. (2024), Kulkarni et al. (2023), and the UVEX Mission website.
@@ 1.3 Challenges for Electromagnetic Follow-up
Numerous thorough searches were performed for the BNS merger GW190425 (Abbott et al., 2020a) (e.g., Antier et al., 2020a; Coughlin et al., 2019b; Gompertz et al., 2020; Hosseinzadeh et al., 2019; Saleem et al., 2020; Song et al., 2019; Paek et al., 2024), as well as the NSBH mergers GW200105 and GW200115 (Abbott et al., 2021a) (e.g., Page et al., 2020; Antier et al., 2020b; Anand et al., 2021; Dichiara et al., 2021) and GW events in O4 (Ahumada et al., 2024). Part of the challenge lies in the broad sky area that any such follow-up endeavor must search. GW detectors do not have high angular resolution to individual sources, and while a network of 3+ detectors allows for significant improvements in this respect, GW localization regions — that is, the area of the sky covered by, e.g., the 90% credible level (C.L.) error region for the GW-inferred localization — are expected to remain large (many 10s or 100s of square degrees) for the immediate future (Petrov et al., 2022; Kiendrebeogo et al., 2023). Tiling a GW localization region spanning such a broad area of the sky is a difficult task for most modern observatories. Adding this difficulty, light curve predictions from KN models for these events are still quite uncertain. This is largely due to the fact that we do not yet understand the dominant mechanism powering the KN at early times. Possibilities include a predominantly nucleosynthesis-powered KN (Li & Paczyński, 1998; Metzger et al., 2010; Hotokezaka & Nakar, 2020; Banerjee et al., 2020; Klion et al., 2021; Banerjee et al., 2024), a nucleosynthesis-powered KN with additional contributions from $\beta$-decay of free neutrons (Metzger et al., 2015; Gottlieb & Loeb, 2020), and a shock-powered KN (Piro & Kollmeier, 2018). Depending on which of these prescriptions is used, the corresponding light-curve predictions can vary by more than 4 mag over the first few hours after merger (see, e.g., Fig. 43 of Kulkarni et al., 2023). When this uncertainty is joined by the intrinsic GW uncertainties on the inclination $\iota$ and luminosity distance $D_{L}$ of the BNS/NSBH merger in question, the depth at which a given telescope needs to observe each field to be able to actually detect a potential KN becomes difficult to estimate. One could, of course, take a highly conservative approach and observe each field at extreme depth, but doing so requires large per-field exposure times.
@@
Time is, however, of the essence. Not only are KNe fast-evolving and fade quickly (as seen for AT 2017gfo in, e.g., Cowperthwaite et al., 2017), but many open questions as to the physical nature of KNe and the origin of heavy elements can only be answered by observing the KN during the rapid evolution of its light curve at early times. As noted above, different underlying mechanisms can cause drastically different KN luminosities in the first few hours. After 5–10 hrs, however, the magnitude difference between (e.g.) shock-powered vs. nucleosynthesis-powered KN models shrinks dramatically, likely to within observational uncertainty of one another (e.g., Fig. 43 of Kulkarni et al. (2023) and surrounding discussion). The exact nature of the early-time evolution of the KN lightcurve is ambiguous, as the first observations of KN AT 2017gfo were performed 15 hrs after the merger (Abbott et al., 2017b; Soares-Santos et al., 2017; Cowperthwaite et al., 2017; Nicholl et al., 2017; Chornock et al., 2017; Margutti et al., 2017; Alexander et al., 2017). Taken together, these considerations both complicate and lend additional impetus to EM-GW follow-up observations, requiring them to be deep, broad, and rapid. This quandary is the driving motivation for employing wide-field survey telescopes in the pursuit of EM follow-up to GW events. These observatories are designed with wide fields of view and the ability to quickly move from pointing to pointing for efficient survey operations. The primary challenge then becomes how one can achieve the necessary depth for successful followup of a given KN counterpart in the short time before it fades without sacrificing coverage of the GW localization region.
@@ 4 Estimated Success Rates
Additional complicating factors (extinction, UV background, etc.) are accounted for within the UVEX ETC. The true apparent magnitude of the counterpart, $m_{\mathrm{KN}}$, is given in terms of the true intrinsic KN absolute magnitude and the actual (simulated) source distance:
@@
The prior bounds are chosen so as to encapsulate the full extent of KN FUV emission models considered in Kulkarni et al. (2023); see Fig. 43 and surrounding discussion therein. We emphasize that this distribution is solely used to estimate the probability of success for the ToO strategies considered in this work as applied to a given observing scenario. The strategies in this work assume a fixed point estimate of the intrinsic KN luminosity with which to compute necessary exposure times and observing schedules. We do not at this juncture have a good understanding of the true distribution of UV KN luminosities, in part because of uncertainty as to the underlying KN UV emission mechanism(s). As such, the choice of a uniform distribution approximates the current broad uncertainty in the absolute magnitudes of the (simulated) KNe in our simulations. In contrast, a probabilistic approach to the ToO selection strategy itself would benefit from assuming a prior that less heavily weights brighter events such as a Gaussian or half-Gaussian distribution. Such an approach is an important aspect of further work on EM-GW follow-up with UVEX (see discussion in §7; Singer et al., 2025).
We can now write an expression for the probability of success in terms of depth, $p_{\mathrm{depth}}(\mathcal{S})$:
@@ 5 Updated Prospects for UVEX EM-GW Follow-up
The prospects presented here for UVEX EM-GW follow-up build on the procedure developed for the UVEX proposal and selection process, bringing it in line with the current post-selection UVEX timeline and mission details — although these remain subject to change — as well as our improved understanding of the BNS rate and mass distribution. As the precise details of the LVK observing run timeline and detector sensitivity in the coming years remain uncertain, we consider two fiducial observing scenarios. The “O5” scenario considers 1.5 years of overlap with a LVK detector network consisting of the LIGO Hanford, LIGO Livingston, Virgo, and KAGRA at their respective projected O5 sensitivities. The “O6” scenario considers the same configuration at the projected sensitivities for O6 and a fiducial 1.5 year overlap. These prospects are estimated by applying our UVEX EM-GW ToO selection criteria as developed for the UVEX proposal to the updated scenarios of Kiendrebeogo et al. (2023) with the GWTC-3 astrophysical BNS rate estimate of $210^{+240}_{-120}\ \mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$ (Abbott et al., 2023). The uvex-followup configuration used for both the O5 and O6 simulations is given in Table 3. These updated estimates retain the UVEX ToO selection strategy developed for the UVEX proposal, and can therefore be taken as a modernization of the quotes presented in Kulkarni et al. (2023).888i.e., 20 ToO triggers for O5 and 35 for O6; the (Kulkarni et al., 2023) quotes did not include the detailed treatment of uncertainty now present within uvex-followup (as described in §3.3). We consider an expanded set of ToO selection strategies in §6.
For the O5 (O6) simulation, with the fiducial observing strategy detailed in Table 3 and an overlap of 1.5 years of concurrent observations, UVEX is estimated to trigger EM-GW ToO observations for $8.9^{+11.1}_{-5.0}$ ($11.3^{+14.0}_{-6.3}$) BNS mergers. The success rate of the fiducial observing strategy as applied to the O5 (O6) scenarios is 76.3% (77.5%), corresponding to an estimated $6.8^{+8.4}_{-3.8}$ ($8.8^{+10.9}_{-4.9}$) KN counterpart detections. The overall distribution of these ToO triggers with respect to their GW distance estimates and sky localizations for O5 and O6 can be found in top panels of Fig. 1. It is expected that the LVK BNS rate measurement will change following O4; to allow the estimates presented in this work to be robust to rate updates in the coming years, we provide a simple tool to convert these quotes to the corresponding quote with a BNS rate of the user’s choice. This tool can be found at https://github.com/criswellalexander/uvex-rate-updater.
@@ 6.2 Coverage Threshold
The percent coverage threshold defines the minimum allowed coverage by UVEX of the GW localization region. It is given as the total integrated probability tiled by the observing plan; e.g., 99% coverage equates to a 99% chance that the counterpart is present within the observed fields. Higher coverage thresholds yield a higher confidence that the UVEX ToO observing plan will include the KN counterpart; however, a stringent threshold may reject events with a high probability of success, but which are missing some small area of the localization region due to sun exclusion, time constraints, etc.. It is worth noting that the threshold defines a minimum level of coverage, and therefore most selected events will in fact encapsulate a greater total probability than given by the threshold. The (quite conservative) fiducial coverage threshold employed for the UVEX proposal process was 99%. We consider two variant strategies — dubbed CT95 and CT90 — in which this threshold is relaxed to 95% and 90%, respectively, and consider the impact on the overall ToO count and success rate. Doing so yields a small increase in ToO triggers with a negligible change in success rate: adopting the CT90 strategy leads to $3.2^{+3.9}_{-1.8}$ and $3.4^{+4.3}_{-2.0}$ additional triggers for the O5 and O6 scenarios, for a total of $12.1^{+15.0}_{-6.8}$ and $14.7^{+18.3}_{-8.3}$ triggers, respectively. The success rate of the CT90 strategy as applied to the O5 scenario is 0.3% higher than that of the fiducial case and is only reduced by 0.2% for the O6 scenario, yielding a total of $9.3^{+11.5}_{-5.2}$ ($11.4^{+14.1}_{-6.4}$) detected counterparts for O5 (O6). The relative ToO trigger counts and success rates for coverage thresholds of 99%, 95%, and 90% can be found in Table 4. An event selection plot for the CT90 strategy can be found in the bottom left panel of Fig. 1. In general, relaxing this requirement results better retention of nearby/well-localized events that fall marginally outside of the UVEX FoR; compare, e.g., the proportion of events lost to FoR exclusion in the O6 simulation with this strategy as opposed to with the fiducial strategy, as seen in Fig. 1.